HETAV PATEL

Grade 11,

An element cystalises in fcc and has edge length 0.56nm. Calculate the density of the element. Molar mass of the element is 40 gmol-1

An element cystalises in fcc and has edge length 0.56nm. Calculate the density of the element. Molar mass of the element is 40 gmol-1

Grade:12

2 Answers

Arun
5 years ago

The density of the crystal is given by the formula:

d = ZM/a^3NA

Given data Z for FCC = 4; M (atomic mass) = 40 g/mol; a (inter-atomic distance) = 0.556 nm or 0.556x10^-9 m, Substituting the values in the formula the density would be

 d = (4*40 g)/(0.556x10^-9m)^3*6.023x10^23

  = 160/(0.1719x10^-27m3) *6.023x10^23

  = 1.546x10^6 g/m3

  d = 1.546 g/cm3

Priya Udupi
2 years ago
d=ZM/a³NA
d=4×40/(5.6×10^-8)³ × 6.023×10²³
d=160/171.87×10–²⁴ × 6.023×10²³
d=160/1035.17×10–¹
d=0.154×10–¹
d=1.54g/mol
🙏🙏🙏🙏🙏